OpenStudy (anonymous):

Pyramid A is a square pyramid with a base side length of 8 inches and a height of 9 inches. Pyramid B has a volume of 2,880 in^3. How many times bigger is the volume of pyramid B than pyramid A? Select one: a. 15 b. 32 c. 54 d. 75

OpenStudy (anonymous):

@Chris911

OpenStudy (anonymous):

@Abmon98

OpenStudy (abmon98):

volume of a pyramid =1/3(b)^2*h so 1/3(64)*9=192 2880/192=15

OpenStudy (anonymous):

Okay so it was literally just plugging in numbers yay